OpenStudy (anonymous):

Age Ticket Price Under 3 years free 3 to 12 years $36 Over 12 years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

(92 - x)36 + (92 - 40) ⋅ 48 = C (40 + x)36 + (92 - (40 + x)) ⋅ 48 = C (92 + x)36 + (92 - (40 + x)) ⋅ 48 = C (40 - x)36 + (92 - 40) ⋅ 48 = C

OpenStudy (anonymous):

Total 92 people. 40 people from ages 3 & Under-12

OpenStudy (anonymous):

Mhmmm

OpenStudy (anonymous):

I believe its either a or d since the number of children under 3 go in for free

OpenStudy (anonymous):

I believe it's D.

OpenStudy (anonymous):

(92-40) represent the amount of student over 12. Since 40 represent 12 years old and less. Then you multiply that by 48 to get the cost. Make sense?

OpenStudy (anonymous):

That it does thanks for the explanation and help